Which strategy is the most appropriate strategy to solve x^2-8x=242?
1
Expert's answer
2019-03-04T13:12:39-0500

The most appropriate strategy to solve x^2-8x=242 is completing the square.


Given quadratic equation x^2-8x=242


Write equation in the form x^2+2ax+a^2=(a+x)^2


Solve for a,


2ax=-8x then we get the value of a is -4


by adding a^2=(-4)^2 to both the sides of given equation,


then we get x^2-8x+(-4)^2 = 242 +(-4)^2


it implies (x-4)^2 = 258


so x-4 = ±√258


therefore x = √258+4 and x = -√258+4


Hence the two zeros of quadratic equation x^2-8x = 242 are x = √258+4, x = -√258 +4.